OpenStudy (anonymous):

always, sometimes, or never true: an integer added to an integer is an integer. an integer multiplied by an integer is an integer. an integer divided by an integer is an integer.

OpenStudy (anonymous):

just did these give me a sec

OpenStudy (anonymous):

1 always 2 always 3 sometimes add sub multiply is always divide is sometimes

OpenStudy (anonymous):

is it the same with polynomials as it is with integers? @missiey

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just did that this morning 100%

OpenStudy (welshfella):

yes thats correct divisiion by integers examples 6 / 3 = 2 - integer result 6 / 4 = 1.5 - non-integer result